e-log length [m]
556 38.70 +/- 0.08 Cavity swinging measurement
556 38.67 +/- 0.03 tape & photo
776 38.640 +/- 0.007 Beam injection, poor PLL, Transmitted DC
this 38.6455 +/- 0.0012 Beam injection, independent PLL, Transmitted DC
Last night, I tried to find the resonance of Yarm by sweeping the frequency of the injection beam.
A strong beat was present at LT_NPRO=48.7856[C_deg], the power coupling of the injection beam was estimated to be 35%.
(Vmax_beat = 1.060[V], Vmin_beat = 0.460[V], Vno_inject = 0.664[V])
The Yarm was locked and the alignment script was executed. The PLL between the PSL beam and the injection beam was
I tried to scan the freq offset (f_PLL) at around 3.88MHz first, then at around 15.52MHz. They are supporsed to be the
first and fourth FSR of the Yarm cavity. The Yarm transmitted power (DC) was observed to find the resonance of the
injection beam. It would have been better to use the RF power, but so far I didnot have the RF PD prepared at the end
transmission. I just used the DC power.
I think I saw the increase of the transmitted power by 10%, at f_PLL = 15.517 +/- 0.003 [MHz]. This corresponds to the
arm cavity length of 38.640 +/- 0.007 [m]. The previous measurement was not so bad!
e-log length [m]
556(2008-Jun-24) 38.70 +/- 0.08 Cavity swinging measurement
556(2008-Jun-24) 38.67 +/- 0.03 tape & photo
This 38.640 +/- 0.007
However, I had difficulties to have more precise measurement mainly because of two reasons:
o The PLL servo is too naive, and the freqency stability of the inj beam is not enough.
The injected beam should have the linewidth (=freq stability) narrower than the cavity linewidth.
o The PLL servo may experience change of the transfer function at around the resonance. The PLL works the other
frequencies. However, close to the resonance, it starts to be unstable.
So the next stuffs we should do is
o Build the PLL just using the incident beams to the ifo, not by the reflected beams.
o Build sophisticated servo to have better frequency stability.
o RF PD at the transmission.
Left the lab with Yarm locked, flipper down, shutter for the NPRO closed.
I tracked the tendency for ezcaPut to fail and sometimes seg-fault in the camera code to a conflict between the camera API and ezca, either on the
network level or the thread level. Since neither are sophisticated enough to provide controls over how they handle these two things, I instead
separated the call to ezcaPut out into a small, separate script (a stripped down ezcawrite), which the camera code calls at the system level. This is a
bit hacky of a solution, but its the only thing that seems to work.
I've developed a transformation based on Euler angles that should be able to take the 4 OSEMs in a picture of the end mirror and use their relative
positions to determine the angle of the camera to the optic. This would allow the position data determined by the fitting software to be converted
from pixels to meaningful lengths, and should aid any servo-ing done on the beams position. I've yet to actually test if the equations work, though.
The servo code needs to have slew rate limiters and maximums/minimums to protect the mirrors written in to it before it can be tested again, but I
have no idea what reasonable values for these limits are.
Joe and I recently scanned the PMC by driving C1:PSL-PMC_RAMP with the trianglewave script over a range of -3.5 to -1.25 (around 50 to 150 volts
to the PZT) and read out C1:PSL-ISS_INMONPD to measure the transmission intensity. This included slightly under 2 FSRs. For slow scans (covering
the range in 150 to 300 s), the peaks were very messy (even with the laser power at 1/6 its normal value), and it was difficult to place where the
actual peak center occurred. For faster sans (covering the range in 30 seconds or so), the peaks were very clean and nearly symmetric, but were
not placed logically (the same peak showed up at two very different values for the PZT voltage in two separate runs). I don't have time to put
together graphs of the scans at the moment; I'll have that up sometime this afternoon.
I was notified by Rob and Rana that there were many measurements of the MC abs length (i.e. modulation
frequencies for the IFO.) between 2002 and now.
So, I dig the new and old e-logs and collected the measured values of the MC length, as shown below.
I checked the presence of the vent for two big steps in the MC length. Each actually has a vent.
The elog said that the tilt of the table was changed at the OMC installation in 2006 Oct.
It is told that the MC mirrors were moved a lot during the vent in 2007 Nov.
o The current modulation freq setting is the highest ever.
o Rob commented that the Marconi may drift in a long time.
o Apparently we need another measurement as we had the big earthquake.
My curiosity is now satified so far.
Local Time 3xFSR[MHz] 5xFSR[MHz] MC round trip[m] Measured by
2002/09/12 33.195400 165.977000 27.09343 Osamu
2002/10/16 33.194871 165.974355 27.09387 Osamu
2003/10/10 33.194929 165.974645 27.09382 Osamu
2004/12/14 33.194609 165.973045 27.09408 Osamu
2005/02/11 33.195123 165.975615 27.09366 Osamu
2005/02/14 33.195152 165.975760 27.09364 Osamu
2006/08/08 33.194700 165.973500 27.09401 Sam
2006/09/07 33.194490 165.972450 27.09418 Sam/Rana
2006/09/08 33.194550 165.972750 27.09413 Sam/Rana
----2006/10 VENT OMC installation
2006/10/26 33.192985 165.964925 27.09541 Kirk/Sam
2006/10/27 33.192955 165.964775 27.09543 Kirk/Sam
2007/01/17 33.192833 165.964165 27.09553 Tobin/Kirk
2007/08/29 33.192120 165.960600 27.09611 Keita/Andrey/Rana
----2007/11 VENT Cleaning of the MC mirrors
2007/11/06 33.195439 165.977195 27.09340 Rob/Tobin
2008/07/29 33.196629 165.983145 27.09243 Rob/Yoichi
27.0924 27.0934 [m]
Alberto and Koji,
Last Friday evening, Koji found that the power adj setting (indicated by ADJ) of the NPRO was somehow set to be
ADJ=-45 and yielded the output power of about 200mW instead of 700mW. This is not good because too small pump power
varies thermal conditions of the crystal such as thermal lensing, thermal gradient, and os on. The ADJ setting and the
crystal temperature had been restored to ADJ=0 and LT=~48deg (nominal of the controller), respectively.
Today we tried the quest of the beating again and the above power setting helped a lot! The beating was immediately
found at LT=48.55deg that is very close to the laser's nominal temp. Also the beating got significantly bigger.
After the alignment adjustment 50%-intenisity modulated signal was obtained. From the power calculation it was
estimated that the power coupling of the injected beam is to be 12%~13%. This not so good yet, but something which we
This time the modulation structure of the PSL beam was clearly observed. I could obtain the beating of the injection
beam with the carrier, the upper/lower sidebands of the 33MHz and 166MHz modulations, and the 2nd order of the
33MHz. They were beautiful as if working with an OSA. Very nice.
In reality, those additional intenisty modulations as well as the residual 33MHz signal from the main IFO are
disturbing for the PLL to be locked at the proper frequency. So, now Alberto is working on a passive LPF with
notch at 33MHz. The design was already done. This allows us to work up to 20MHz and at the same time, provides
60dB attenuation at 33MHz (in principle). Very cool.
Koji, on the other hand, continued to work with the PLL servo with some ready-made passive filters. Owing to the
fillters, the error signal was cleaner and the PLL was locked at the proper frequency. The PLL setup is as attatched.
Sideband rejection filter will be replaced to Alberto's one. The photo is the display of the RF spectrum analyzer with
beat locked at 8MHz.
So the next step, we try to find the resonances of the arm cavity with the injection beam once the IFO comes back.
At the last of the experiment "Last autoalignment" was restored, the flipper for the
inj beam was down, and the shutter for the NPRO was closed.
1) Continue mode matching into the PMC. Its transmission now is around the same as the
2) Put a UHV foil covered lead brick onto the PMC to quiet it down.
3) Characterize the PMC loop and retune the body notch for the new body.
4) Tweak the MZ alignment to minimize the RFAM. We can use StochMon to do this as
long as we have the MC WFS turned off or we can put in a flipper to take the
beam before the MC and send it to the StochMon RFPD.
5) Re-align onto the ISS.
6) Install irises around the periscope for the beam. The old iris there is way off.
7) Fix PSL ANG and center both POS and ANG.
Alberto and Koji
o We worked for the abs length measurement setup on Thursday night.
o At the last of the work Koji left the 40m lab at 2AM. "Last autoalignment" was restored. The flipper for the
inj beam was down. The shutter for the NPRO was closed.
o The alignment of the injection beam (NPRO) was re-adjusted.
o The laser crystal temp (LT) of the NPRO was scanned.
o After a long struggle the beat was found at about LT=61deg(!). I think this is almost at the maximum temp
for the NPRO. Note that this is not the diode temp, and therefore it will not damage the laser as far as the
TEC for the crystal works.
o Only the X arm was aligned.
o The alignment of the injection beam was adjusted such that the beating amplitude got maximum.
o At the faraday of the NPRO, we had 2.4V_DC and 1.8V_DC with and without the inj beam, respectively. The
beating amplitude was 200mVpp (at around 2.4V).
o With a simple calculation, the mode overlapping of tghe injection beam is only 0.0023. Ahhh. It is too weak.
In the modematching or something must be wrong.
o The position of the mode matching lens was tweaked a little. It did not help to increase the beat ampitude.
Even worse. (The lens was restored and the values above was obatined with the latest setting.)
o Then tried to build a PLL. It locks easily.
- Put the beat signal into the mixer RF input.
- Connect 10dBm @1MHz-10MHz from the marconi oscillator to the LO input. The supposed nominal LO level was
not checked so far. Just used 10dBm.
- The IF output was connected to an SR560 with 10Hz LPF (6dB/oct) with G=500 or so.We don't need to care
about the sign.
- The output of the SR560 was connected to FAST PZT input of the NPRO.
o The problem was that there was strong intermodulations because of 33MHz. No LPFwas used before the mixer.
Because of this spourious modulations, the PLL servo locks at the local zero crossings. These will be solved
o Eventually left the 40m lab at 2AM. "Last autoalignment" was restored. The flipper for the inj beam was
down. The shutter for the NPRO was closed.